Find X in log6 x= ‐2


1
Expert's answer
2021-12-06T16:32:15-0500

Solution.

For equation "\log_ax=b" we find solution as "x=a^b."

So, we have

"\log_6x=-2""x=6^{-2}\newline 
x=(\frac{1}{6})^2""x=\frac{1}{36}"

Answer. "x=\frac{1}{36}."